### v1

# GSQA: An End-to-End Model for Generative Spoken Question Answering

Abstract
In recent advancements in spoken question answering (QA), end-to-end models have made significant strides. However, previous research has primarily focused on extractive span selection. While this extractive-based approach is effective when answers are present directly within the input, it falls short in addressing abstractive questions, where answers are not directly extracted but inferred from the given information. To bridge this gap, we introduce the first end-to-end Generative Spoken Question Answering (GSQA) model that empowers the system to engage in abstractive reasoning. The challenge in training our GSQA model lies in the absence of a spoken abstractive QA dataset. We propose using text models for initialization and leveraging the extractive QA dataset to transfer knowledge from the text generative model to the spoken generative model. Experimental results indicate that our model surpasses the previous extractive model by 3% on extractive QA datasets. Furthermore, the GSQA model has only been fine-tuned on the spoken extractive QA dataset. Despite not having seen any spoken abstractive QA data, it can still closely match the performance of the cascade model. In conclusion, our GSQA model shows the potential to generalize to a broad spectrum of questions, thus further expanding spoken question answering capabilities of abstractive QA. Our code is available at https://voidful.github.io/GSQA
Index Terms—  Spoken Question Answering, Textless NLP, Generative Language models, Transfer learning
1 Introduction
Question Answering (QA) tasks have consistently emerged as one of the foundational challenges [1, 2, 3, 4]. As a touchstone for measuring machine comprehension of textual content, QA presents unique complexities that have long captivated researchers. If answers are directly in the text and just need to be found, we categorize it as extractive QA. In contrast, when the answers are not explicitly present, needing a deeper dive involving inference and synthesis to craft a response, we refer it to as abstractive QA. With the rise of smart speakers and voice assistants, an emerging challenge in QA is Spoken Question Answering (SQA) to facilitate interactions between humans and these assistants. SQA requires both the questions and answers to be vocalized. Unlike text, speech encompasses richer details such as the identity of the speaker, their tone, and more, which introduces added intricacies to the SQA tasks.
An intuitive approach to tackle this challenge is the cascade method. Here, an Automatic Speech Recognition (ASR) system first transcribes the input speech of both questions and the passages, and then feeds transcriptions into a textual QA model that predicts the corresponding answers. Last, answers are transformed back into speech through a Text-To-Speech (TTS) model. The cascade method’s primary advantage is its ability to harness the power of text-based QA models trained on vast datasets, ensuring it can tackle both extractive and abstractive QA tasks. However, the cascade model has its drawbacks. The primary issue stems from error propagation, since the text-based QA model isn’t typically trained on ASR-erroneous data, the presence of ASR errors can notably undermine the performance of cascade models.
Fig. 1: GSQA compared to other baselines: The Cascade Method accommodates both abstractive and extractive QA but risks error propagation. DUAL is an end-to-end textless approach, exclusive to extractive QA. GSQA is a textless, end-to-end generative method, capable of handling both extractive and abstractive QA.
Recently, end-to-end, textless approaches (e.g., DUAL[5]) have emerged as a solution to the cascade method’s limitations. These works [5, 6, 7] rely on the representations of acoustic models, like HuBERT[8], and apply K-means clustering to transform speech into discrete units. These units, encapsulating both questions and passages, are fed into the model that predicts the span of the answer in the passage using start and end positions. By adopting this textless approach, DUAL effectively sidesteps the error propagation from ASR systems [9, 10, 11]. It improves the model robustness in noisy speech. However, there’s a notable limitation to the DUAL model: it is not able to handle abstractive QA scenarios as its design to only output start and end positions within the input.
As Fig.1 shows the different of our purpose method, cascade method, and DUAL, we combine the advantages of both the cascade model and the end-to-end textless approach to introduce the first-ever textless end-to-end model capable of handling both extractive and abstractive QA. Mirroring the DUAL’s methodology, we utilize HuBERT to convert the input speech (question and passage) into discrete units[6]. A sequence-to-sequence model then processes these units to generate the answer in discrete units. Motivated by the strengths and limitations of approaches above, we propose GSQA (Generative Spoken Question Answering), the first textless end-to-end model capable of handling both extractive and abstractive QA.
Furthermore, given the current absence of a spoken abstractive QA corpus, we begin by initializing GSQA with a generative Text QA (TQA) model. This TQA model has been trained on several extractive and abstractive datasets using a generative approach. Subsequently, we fine-tune the TQA model on NMSQA, a dataset dedicated to extractive spoken QA. GSQA primary objective is to impart the Natural Language Understanding prowess of the TQA model to GSQA. We also use Microsoft’s Azure TTS service to create a synthesized test set for evaluation. It’s noteworthy that, during the transfer learning phase, GSQA exhibited a significant enhancement, registering an 18% relative gain in the BLEU1 score in a zero-shot abstractive spoken QA setting.
The contributions of this paper are summarized below:
  1. 1.
    The introduction of spoken generative question answering model GSQA, the first end-to-end textless generative model.
  2. 2.
    The establishment of a method where pre-training on textual QA, followed by fine-tuning on extractive spoken QA, leads to zero-shot abstractive spoken QA.
  3. 3.
    Demonstrating that our generative QA model has competitive performance against textless extractive models.
2 Related Work
Most existing SQA methods rely on Automatic Speech Recognition (ASR) transcripts and then use a large language model (LLM) to fulfill the target of question-answering tasks. However, cascade models are strongly influenced by the ASR errors when an ASR system [12, 13] is not so robust that would pass wrong documents to LLM, and that causes LLM to predict wrong answers. On the other hand, a robust ASR system, like Whisper-large [13], which has 1550M parameters, can achieve a low word error rate, and can pass accurate transcriptions of documents to LLM to execute QA tasks well. However, though cascade models can achieve quite surprising accuracy on QA tasks when leveraging such an awesome ASR system, the large number of models’ parameters would take longer inference time.
Extractive SQA is to find the answer from a spoken document given a question. DUAL [5] is designed to predict the start and end index of answer spans. However, the limitation of DUAL is that it can only predict the answer span of the input audio, which means if answers do not exist in given spoken documents, it becomes useless.
TWIST [14] is an approach for solving the limitation of SpeechLM [15], which is the lack of semantic understanding that might lead to off-topic or wrong responses. However, though TWIST finds a better initialization point for SpeechLM, the experiment results of TWIST still do not show it has major semantic knowledge transfer.
3 Methodology
3.1 Overview & Formulation
In the realm of Question Answering (QA), we divide the problem into three components: the Question (Q), the Passage (P), and the Answer (A). These entities can manifest in two formats: textual data ($Q_{t}$, $P_{t}$, $A_{t}$) and speech discrete unit data ($Q_{u}$, $P_{u}$, $A_{u}$).
Our method lies in the ambition to create a textless generative end-to-end QA model. Initially, we follow [6, 7] to quantize speech into discrete units. We then utilize a discrete unit-based sequence generation model to craft the answer units, which would be subsequently transformed into speech via a Unit-based HiFi-GAN vocoder [16]. To boost the logical understanding capabilities of the discrete unit-based sequence generation model, we initialize this generative spoken QA model using weights from a textual model [14]. Fig.2 shows the entire model architecture and training pipeline.
Fig. 2: Left: The process of discrete unit quantization from synthesis data. Right: Model Training Procedure: A depiction of the transition from textual QA pre-training to spoken QA fine-tuning.
3.2 Speech Quantization
The HuBERT[8] model serves as a formidable speech representation model. It collects a general representation by undertaking mask segment prediction training on its intermediate representation using k-means clustering. During this, every 20ms frame is clustered into one of the K categories. We obtain the k-means centroid id as the discrete representation of the frame, effectively transforming speech signals into discrete units. Following the methods presented in Lee et al. [6], we employed the HuBERT-base model and a 100 cluster number setting to achieve this transformation. Furthermore, to accelerate the process, we adopt the run-length encoding strategy to collapse consecutive sequences of identical units into a singular unit, thereby shortening the sequence length.
3.3 Text QA Model Pretraining
Our generative spoken QA model’s training data is relatively scarce, which could hamper the model’s performance. To tackle this issue, we suggest leveraging a pretrained textual model to amplify its semantic understanding. Moreover, this textual model is initially trained as a generative QA model before being transferred to a spoken QA model. This approach ensures that even if our spoken QA data is limited, the model can learn to transfer from text to speech units, allowing the capabilities of textual QA to be transferred onto the unit.
When converting speech to units, and even after deduplicating tokens, the resulting speech sequence remains longer than the text. Many generative textual models[17] can not handle such long input sequences. Therefore, when choosing our initialization textual model, it had to fulfill two criteria: 1) Be a generative model, and 2) Handle sequences exceeding 4096 in length[18, 19, 20].
In the end, we choose LongT5[18]. LongT5 extends the original T5 encoder, integrating global-local attention sparsity patterns to manage extended inputs. It marries the efficient attention mechanism of ETC with the summarization-centric pre-training objective of PEGASUS, resulting in exceptional performance. Upon this foundation, we further pretrain LongT5 on textual QA datasets.
The primary ambition during the pre-training phase is to develop a model capable of accepting a question $Q_{t}$ and passage $P_{t}$ in textual form and predicting the corresponding answer $A_{t}$. This training procedure is devised to equip our model with the essential capabilities needed for the ensuing stages of our method.
3.4 GSQA: Generative Spoken QA Model
Fig.2 illustrates our GSQA training procedure. Initially, we train the LongT5 on Text QA to produce the TQA LongT5 model, which is capable of text-level generative QA. This model is then fine-tuned on the Extractive Spoken QA dataset, NMSQA. We converted the NMSQA into a generative QA format, and it will output the answer directly in discrete unit form instead of the answer’s start and end position from extractive QA. Prior to this, both the input and output audio go through quantization to be converted into discrete units. We also employ a vocoder, based on the HiFi-GAN vocoder from [21], to convert these discrete units back to speech. This vocoder includes a duration prediction module for deduplicated unit synthesis. Since TQA-LongT5 doesnt́ recognize discrete units, we introduce them as new tokens. Their embeddings are initialized by random sampling from other text token embeddings, ensuring proper tokenization and a seamless transition from text to speech.
NMSQA-dev NMSQA-test Spoken-NarrativeQA-test Model Num. Param. F1 Score EM Score F1 Score EM Score BLEU1 ROUGE-L Cascade model (Upper bound) 1025M 49.1 32.0 47.3 30.4 13.5 19.9 DUAL 452M 39.4 21.9 33.6 21.2 - - Unit-LongT5 312M 25.5 12.6 20.1 9.4 6.8 10.4 Unit-LongT5-TQA (Proposed) 312M 41.8 24.9 36.0 24.0 8.0 11.8
Table 1: The table presents the Speech to Speech Evaluation on our models and the number of parameters for each method. For extractive QA datasets, we utilized F1 and exact match (EM) as evaluation metrics. On the other hand, for generative QA datasets, we employed BLEU1[22] and Rouge-L[23] for assessment. The cascade model in this table includes an ASR system and a LM, which is Whisper-medium.en [13] and LongT5-TQA[18], respectively. DUAL[5] is combined a HuBERT-base[8], K-means model for 128 clusters, and a Longformer[24].
4 Experimental Setup
4.1 Dataset
We use NMSQA dataset [5] as extractive SQA dataset. The NMSQA is derived from SQuAD v1.1, consists of Wikipedia paragraphs with human-written questions. While its train and dev sets feature synthesized speech using Amazon Polly, the test set contains audio from 60 human speakers (30 males and 30 females). In addition, we propose a new spoken abstractive SQA evaluation set, named Spoken-NarrativeQA. This dataset is built upon NarrativeQA test set, a abstractive text QA dataset for reading comprehension from books and movie scripts with complex narrative understanding. Note that many answers in this dataset are not directly extractable from passages. We transform the text to speech using TTS system and specifically select test data where answers are absent in the passage. There are 1626 testing samples in the Spoken-NarrativeQA dataset. For TQA training data, We utilized four extractive QA datasets: SQuAD [25], MultiRC [26], NewsQA [27], and Drop [28], along with one abstractive dataset, NarrativeQA [29], to train LM, and use the trained LM as foundation model of GSQA.
[…]
Table 2: The respective number of parameters in our experiment. Note that E2E method, like DUAL and our work, does not need ASR model but need Unit-Extraction model to generate units, and the number after the name of that is the number of clusters.
NMSQA-dev NMSQA-test Spoken-NarrativeQA-test 8.1 14.8 6.3
Table 3: Word Error Rate (WER) of the pretrained Whisper-medium.en ASR system on different datasets for Spoken QA tasks.
4.2 Cascaded Pipeline
The conventional approach for SQA tasks typically follows a cascaded pipeline, as illustrated in Figure 1 middle. The passage and question are first transcribed using ASR, and then a text-based QA model generates the textual answer. To generate the final spoken answer, we utilize the Azure TTS service to generate the spoken answer. This methodology serves as our baseline and is referred to as the Cascade Model. More specifically, our ASR model is using Whisper-medium.en [13], which has the state-of-the-art performance with comparable parameter size to DUAL and ours model. For the QA component, we employ the fine-tuned LongT5-TQA model, as mentioned in Section 3.3. The ASR performance of datasets is in Table 3.
4.3 Performance Evaluation
Since all our models ultimately output as a waveform, what we focus on in QA is whether the content of this waveform contains the answer. We then use an ASR model to convert the generated waveform back to text, allowing us to use text-based metrics to evaluate the content of the model. To reduce errors from the ASR during evaluation, we utilize the state-of-the-art ASR model, whisper-large-v2, during performance evaluation. We then compare the transcriptions of the outputs from whisper-large-v2 with the text ground truth. We calculate the F1-score and Exact-matched (EM) scores for extractive QA datasets. The BLEU score [22] and the Rouge score [23] are used for generative QA datasets. Lastly, the experimental results are reported in Table 1.
4.4 Implementation details
As the previous section mentioned, we pretrain LongT5 on the 5 TQA datasets for 13 epochs, and the learning rate is set to 0.0005, and weight decay is 0.01. On the other hand, when we fine-tune the pre-trained LongT5 on a unit-based dataset, NMSQA, for 25 epochs, the learning rate is set to 0.0003, and weight decay is 0.001. Last, when we inference our model on abstractive test set(e.g., Spoken-NarrativeQA-test), to encourage GSQA output longer length, we use beam search algorithm for decoding with a number of beam 5, and set length penalty to 2.
5 Results
As illustrated in Table 1, we evaluate on both extractive QA dataset NMSQA-dev, NMSQA-test, and abstractive QA dataset Spoken-Narrative QA. The following discussions provide insights into our main experiment results from different aspects.
5.1 Question Answering Performance
Among the models evaluated, the proposed Unit-LongT5-TQA exhibits remarkable performance. On the NMSQA-dev set, it achieves an F1 score of 41.8% and an EM score of 24.9%. Its performance remains consistent on the NMSQA-test dataset, registering scores of 36.0% and 24.0% for F1 and EM, respectively. Furthermore, this model outperforms the others on the generative Spoken-NarrativeQA-test set, securing a BLEU1 score of 8.0% and a ROUGE-L score of 11.8%.
5.2 The effectiveness of TQA pretraining
From Table1, we can observe that Unit-LongT5-TQA has huge improvements on both NMSQA-dev and NMSQA-test, which gets 16 F1-scores gain than Unit-LongT5, and gets two times than it on EM scores. It shows that using pretrained LM on TQA tasks as the initialization brings significant advantages for LM to learn speech information. On the other hand, in Spoken-NarrativeQA-test, an abstractive spoken QA dataset, Unit-LongT5-TQA also outperforms Unit-LongT5 on both BLEU and ROUGE scores, which gets 1.2 and 1.4 increase, respectively. In our showcase, we have observed that TQA pretraining can help the model generate more fluent sentences. The following results are from our test set. Without TQA pretraining, our prediction is “live the life of any,” where the sentence is truncated in the middle, resulting in an incoherent meaning. However, after TQA training, the model can output fluent sentences.
Question What did Tancred’s destiny seem to be? Answer To live the life of a normal member of the British ruling class. Unit-LongT5 live the life of any Unit-LongT5-TQA live the life of any conventional member of the British ruling class
Table 4: Comparison of Results
5.3 Performance at different WERs
Fig. 3: We sample the 8000 data within NMSQA-dev to verify the impact on cascade model under different Word Error Rates(WERs). Compared to cascade model, Unit-LongT5-TQA is more stable when higher WER leads to cascade model’s poor performance.
From the Fig.3, we can observe that Cascade model is quite sensitive to the impact of ASR WER. When ASR WER is low, we can see that cascade model performs well, which can get  50 F1-score. However, when the ASR WER increases, the drop of cascade model’s performance is obvious, dropping from 50 to 40 F1-score. On the other hand, from Table1, though our model (Unit-LongT5-Unit) performance is not as good as cascade model’s on average, we can see that our model can be more stable when ASR WER increases, even has better performance on high WER than cascade model.
5.4 Models parameters
For a fair comparison, we strive to ensure the number of parameters across all models remains relatively consistent. Upon examination, the cascade approach is found to have a larger number of parameters. This is primarily because it requires stacking various models together. Since the performance of each model should not be compromised, opting for smaller models is not feasible, leading to a larger overall parameter size. In contrast, end-to-end models rely on a single model to accomplish the task, resulting in a reduced number of parameters. We observed that for generative QA, there isn’t a stringent requirement for an exceptionally powerful speech-to-unit model. It’s sufficient to encode the content’s information into discrete units. Instead, a more potent QA model is needed to infer the answers to the questions. Detailed parameter counts are presented in Table 2. It is evident that our model has the fewest parameters among all methods. Notably, our model outperforms the others across various metrics.
6 Conclusion
In conclusion, we introduced the first end-to-end Generative Spoken Question Answering (GSQA) model, bridging the gap between extractive and abstractive QA tasks. By combining the advantages of cascade and textless models, our GSQA model demonstrates superior performance on extractive QA datasets, outperforming extractive spoken model. Furthermore, in the challenging abstractive zero-shot domain, our model exhibits competitive capabilities. We also highlighted the importance of pre-training the LM on textual data to enhance the unit-to-unit model’s semantic understanding. Future research may focus on improving the model’s performance with human narration and exploring strategies to enhance generalization across diverse spoken QA tasks.